OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate 9.2^2 as follows: Let f(x)=x^2 and find the equation of the tangent line to f(x) at x=9. Using this, find your approximation for 9.2^2

OpenStudy (anonymous):

cheat and use a calculator

OpenStudy (anonymous):

F'(x)=2x

OpenStudy (anonymous):

dy=2x dx

OpenStudy (anonymous):

i dont even know where to start or how to cheat with a calculator

OpenStudy (anonymous):

2(9)(.2) 3.6

OpenStudy (anonymous):

or find the equation of the line tangent to the graph of \[y=x^2\] at the point (9,81). the slope is 18 so the line is \[y=18(x-9)+81\] now replace x by 9.2 and you get it

OpenStudy (anonymous):

ok we go slow

OpenStudy (anonymous):

81+3.6=84.6

OpenStudy (anonymous):

you want the equation of the line tangent to the graph of \[y=x^2\] at the point (9,81)

OpenStudy (anonymous):

to fine the equation for the line you need two things: a point and a slope. you already have the point. to find the slope, take the derivative of \[x^2\] which is \[2x\] and replace x by 9 to get the slope of 18

OpenStudy (anonymous):

using the "point - slope" formula you know the equation for the line tangent to \[y=x^2\] at (9,81) is \[y-81=18(x-9)\] \[y=18(x-9)+81\]

OpenStudy (anonymous):

now replace x by 9.2 to get \[y=18(9-9.2)+81\] \[y=18\times .2+81\] \[y=3.6+81\] \[y=84.6\]

OpenStudy (anonymous):

imranmeah way is much snappier, but maybe not so obvious. either way

OpenStudy (anonymous):

btw \[9.2^2=84.64\] so this is just an approximation. that is what i meant when i said cheat and use a calculator. just square 9.2

OpenStudy (anonymous):

ohh thanks a bunch man this was a great help

OpenStudy (anonymous):

yw